OpenStudy (anonymous):

Compare the expressions using <, >, or =. Show all of your work. (14 – 6) ÷ (6 – 2) _____ 4 • (10 – 8) – 2 • 2

OpenStudy (anonymous):

(14 - 6) ÷ (6 - 2) = (8) ÷ (4) = 2 4 • (10 - 8) - 2 • 2 = 4 • (2) - 4 =8 - 4 = 4 As 2 < 4 (14 – 6) ÷ (6 – 2) < 4 • (10 – 8) – 2 • 2

OpenStudy (anonymous):

Thank you!